OpenStudy (anonymous):

can some one show me the steps for ^-3 multiplied by ^-128

Directrix (directrix):

Is this the problem: x^-3 multiplied by x^-128 ? I do not see a variable in the problem you wrote.

OpenStudy (anonymous):

\[\sqrt{-3} \times \sqrt{-128}\]

Directrix (directrix):

Okay. ^ means square root.

OpenStudy (anonymous):

yes.

Directrix (directrix):

√(-3) * √(-128) = √(-1) * √3 * √(-1) * √128 =

Directrix (directrix):

i √3 * i * √128 where i = √(-1).

Directrix (directrix):

= i^2 * √3 * √128 = (-1) * √3 * √128 Questions so far?

OpenStudy (anonymous):

nope

Directrix (directrix):

= (-1) * √3 * √(4*4*2*2*2) = (-1) √3 * (4 * 2)√(2) = - √3 * 8√2

Directrix (directrix):

= -8√6

Directrix (directrix):

Questions?

OpenStudy (anonymous):

so the answer doesn't have and i in it meaning. I got \[8i \sqrt{6}\]

Directrix (directrix):

With complex numbers that are not real, you have to get the negative out of the square root before you multiply the radicands.

Directrix (directrix):

i^2 = -1

OpenStudy (anonymous):

i c. thanks so much.

Directrix (directrix):

Glad to help.